# Resolving the Identity Crisis in Text-to-Image Generation

### v2

@@ 3.4 DisCo Algorithm
We provide the complete Pseudocode for DisCo finetuning in the Supplementary Material (Section A.3). For each update, we sample $n\sim p_{t}(\cdot)$, a prompt $c\in\mathcal{P}_{n}$, generate a group $G$ of $M$ trajectories under the SDE policy, detect faces and compute embeddings, evaluate rewards via Eqs. 610, compute advantages via (3), and update $\theta$ with (4). In the following Section, we discuss the Results of training using DisCo.
4 Experiment
4.1 Experimental Setup
4.1.1 Datasets
Training Data.
DiverseHumans.
MultiHuman-TestBench.
We further evaluate on MultiHuman-TestBench (MHTB), an established recent benchmark introduced at NeurIPS 2025 for multi-human generation. MHTB provides comparison protocols on general multi-human generation capabilities without specific emphasis on identity diversity, and extend the scope of images to people performing simple and complex actions, complementing our DiverseHumans evaluation. Additional details are in the Supplementary Material (Section B).
@@ 4.1.2 Models
We compare against several baseline models including Nanobanana [5], SD3.5 [8], FLUX [10], Krea [11], HiDream-Full [3], Qwen-Image [23], OmniGen2 [24], DreamO [19] and GPT-Image [20]. We fine-tune two open source models, FLUX-Dev(generalist) and Krea-Dev(specialist), using our DisCO framework to allow a direct performance comparison with their baseline counterparts. All implementation details and hyperparameters are provided in the Supplementary Material (Section D).
@@ 4.1.3 Metrics
To evaluate the performance of our model against the baseline, we report three key metrics: Count Accuracy measures the percentage of generated images that contain the exact number of individuals specified in the prompt. Unique Face Accuracy (UFA) quantifies the proportion of images in which all depicted individuals correspond to visually distinct identities, ensuring no duplicates within a single image. Global Identity Spread (GIS) is a global metric and assesses identity diversity across a dataset. by computing the ratio of total unique identities to the total prompted identities, in the testset. It indicates how effectively the model avoids repeating the same identities across different images. HPSv2 assesses image quality and prompt/image alignment. We measure the MLLM Action scores for alignment with textual actions as proposed in MultiHuman-TestBench. See the Supplementary Material (Section C) for the full mathematical details.
@@ 4.3 Ablation Study
Table 2 ablates individual contributions of each DisCo component. This analysis is conducted on the Krea-Dev baseline, which proved more challenging to converge compared to Flux-Dev.
Intra-image diversity dramatically improves unique face accuracy but leaves Global Identity Spread limited, as duplicate identities simply spread across different images rather than being eliminated. Adding group-wise diversity addresses this by enforcing diversity across the entire generation group, substantially improving cross-image identity variation.
Count accuracy drops when applying only group-wise rewards due to reward hacking. The model exploits generating fewer people as an easier optimization target. The count control component provides essential regularization, recovering count performance while maintaining identity diversity. However, this introduces perceptual quality issues including unnatural “grid” arrangements of faces that technically satisfy requirements but appear artificial. HPS quality control effectively mitigates these artifacts by penalizing obvious visual anomalies.
The curriculum learning component delivers substantial improvements. Since Flux-Krea is not a generalist model, training convergence proved challenging without proper task decomposition. Curriculum learning addresses this by progressing from simple to complex scenarios, enabling the specialized model to learn the difficult multi-human generation task incrementally.
As evident from the scores, each component contributes meaningfully to the final performance, with the complete framework achieving optimal results across all metrics despite the challenging baseline characteristics.
@@ 5 Conclusion
Current text-to-image models face a persistent identity crisis in multi-human generation: duplicating faces, merging identities, and miscounting individuals. Prior work focuses on intra-image realism, overlooking identity diversity across samples, which is critical for scalability. We propose DisCo, an RL framework with a compositional reward that enforces intra-image uniqueness and introduces a novel group-relative term to discourage cross-sample identity repetition. Combined with GRPO and a single-stage curriculum, DisCo improves identity diversity and count accuracy without extra annotations or quality loss. Empirically, DisCo achieves substantial gains in identity diversity and count accuracy across diverse prompts and model backbones. On DiverseHumans, DisCo(Flux) reaches 98.6% Unique Face Accuracy, eliminating identity duplication as compared to 48.2% for baseline Flux and 57.2% for proprietary Gemini-Nanobanana. Similar improvements are observed on MultiHuman-TestBench, with DisCo(Flux) achieving 94.3% versus 56.5% baseline. Crucially, these identity-focused optimizations enhance rather than compromise perceptual quality, establishing a new paradigm for multi-human generation that resolves the identity collapse and sets a benchmark for cross-sample diversity in generative models.
@@ E.1.3 Intra-Image Diversity Aggregation Function Analysis
E.2 Final Run Reward Curves
@@ E.2.1 Computational Analysis
E.3 Qualitative Results
E.3.1 Visualizing Global Identity Spread
@@
Figure E.2: DisCo v/s Flux-Dev As observed in this Figure, we visualize three prompts of people containing the same ethnicity, over six consistent seeds for DisCo and Flux. As observed, Flux results not only generate overlapping identites in the same image, but generate similar looking people across the dataset. However, DisCo finetuning pushes the faces further from each other.
E.3.2 Results on Flux-Krea
@@
E.3.3 Visual Effects of Count and HPS Reward Components
@@
Appendix F Limitations and Future Work
Our approach relies on face detection and face-embedding similarity; as such, failure cases can arise under heavy occlusion, extreme poses, partial profiles, or when faces are very small. Future directions for this line of work include integrating body/appearance cues beyond faces (e.g., re-identification or whole-body embeddings), extending DisCo to videos with spatiotemporal identity consistency, extending disco to other (diverse in nature) concepts such as animals, learning adaptive curricula, and exploring human-in-the-loop or active reward shaping. Finally, we aim to study fairness and demographic balance more explicitly, and to evaluate robustness to higher person counts.
Ethics Statement

### v3

@@ 3.4 DisCo Algorithm
We provide the complete Pseudocode for DisCo finetuning in the Appendix A.3. For each update, we sample $n\sim p_{t}(\cdot)$, a prompt $c\in\mathcal{P}_{n}$, generate a group $G$ of $M$ trajectories under the SDE policy, detect faces and compute embeddings, evaluate rewards via Eqs. 48, compute advantages via (1), and update $\theta$ with (2). In the following Section, we discuss the Results of training using DisCo.
@@
4 Experiment
4.1 Experimental Setup
4.1.1 Datasets
Training Data.
DiverseHumans.
MultiHuman-TestBench.
We further evaluate on MultiHuman-TestBench (MHTB), an established recent benchmark introduced at NeurIPS 2025 for multi-human generation. MHTB provides comparison protocols on general multi-human generation capabilities without specific emphasis on identity diversity, and extend the scope of images to people performing simple and complex actions, complementing our DiverseHumans evaluation. Additional details are in the Appendix B.
@@ 4.1.2 Models
We compare against several baseline models including Nanobanana [10], SD3.5 [9], FLUX [14], Krea [15], HiDream-Full [5], Qwen-Image [28], OmniGen2 [29], DreamO [23] and GPT-Image [24]. We fine-tune two open source models, FLUX-Dev(generalist) and Krea-Dev(specialist), using our DisCO framework to allow a direct performance comparison with their baseline counterparts. All implementation details and hyperparameters are provided in the Appendix D.
@@ 4.1.3 Metrics
To evaluate the performance of our model against the baseline, we report three key metrics: Count Accuracy measures the percentage of generated images that contain the exact number of individuals specified in the prompt. Unique Face Accuracy (UFA) quantifies the proportion of images in which all depicted individuals correspond to visually distinct identities, ensuring no duplicates within a single image. Global Identity Spread (GIS) is a global metric and assesses identity diversity across a dataset. by computing the ratio of total unique identities to the total prompted identities, in the testset. It indicates how effectively the model avoids repeating the same identities across different images. HPSv2 assesses image quality and prompt/image alignment. We measure the MLLM Action scores for alignment with textual actions as proposed in MultiHuman-TestBench. See Appendix C for the full mathematical details.
@@ 4.2.1 Quantitative Scores
4.2.2 Qualitative Results
4.2.3 Generalizability and Compositional Control
Beyond facial identity, we demonstrate in Appendix F.2 that DisCo’s reward framework generalizes to other visual attributes. Replacing the ArcFace reward with a CLIP-based [26] clothing similarity signal yields improved clothing diversity across generated individuals. Notably, training on facial diversity alone already improves clothing and body-type variation(measured using Qwen3-VL [1]) as a byproduct. This suggests that the model learns a broader notion of inter-person distinctiveness. On GenEval2 [12], DisCo improves over baseline Flux across object, counting, positional, and verb understanding attributes, demonstrating that diversity-focused RL fine-tuning does not hurt, and in fact strengthens general text-to-image capabilities (Appendix F.3). Finally, despite no explicit training for per-person attribute control, DisCo faithfully follows fine-grained compositional prompts specifying clothing style and hairstyle per individual, as shown in Appendix F.4.
@@ 4.3 Ablation Study
Table 2 ablates individual contributions of each DisCo component. This analysis is conducted on the Krea-Dev baseline, which proved more challenging to converge compared to Flux-Dev. Intra-image diversity dramatically improves unique face accuracy but leaves Global Identity Spread limited, as duplicate identities simply spread across different images rather than being eliminated. Adding group-wise diversity addresses this by enforcing diversity across the entire generation group, substantially improving cross-image identity variation. Count accuracy drops when applying only group-wise rewards due to reward hacking. The model exploits generating fewer people as an easier optimization target. The count control component provides essential regularization, recovering count performance while maintaining identity diversity. However, this introduces perceptual quality issues including unnatural “grid” arrangements of faces that technically satisfy requirements but appear artificial. HPS quality control effectively mitigates these artifacts by penalizing obvious visual anomalies. The curriculum learning component delivers substantial improvements. Since Flux-Krea is not a generalist model, training convergence proved challenging without proper task decomposition. Curriculum learning addresses this by progressing from simple to complex scenarios, enabling the specialized model to learn the difficult multi-human generation task incrementally. As evident from the scores, each component contributes meaningfully to the final performance, with the complete framework achieving optimal results across all metrics despite the challenging baseline characteristics. Additional ablation studies, such as varying the curriculum bounds, grid search on reward weights and varying the intra-image aggregation function can be found in Appendix E. We also provide latency trade-offs as well as a user preference study for our proposed metrics.
@@ 5 Conclusion
In this paper, we uncover and resolve the identity crisis of text-to-image models: when asked to generate a crowd, they consistently reuse the same face. We propose DisCo, fine-tuning flow-matching models via GRPO with a compositional reward built around a novel finding: optimizing diversity within an image is not enough, as duplicates simply migrate across samples instead. Our group-relative reward catches this cross-sample collapse, complemented by count accuracy, perceptual quality, and text alignment terms to keep generations grounded. Combined with a single-stage curriculum, DisCo turns a model that copies faces into one that generates distinct individuals at scale, surpassing proprietary models on identity diversity without sacrificing quality or requiring extra annotations. Beyond faces, our rewards generalize to clothing and body-type diversity, and we demonstrate that RL fine-tuning retains general compositional capability. Future directions include extending identity-aware RL to personalized generation, video generation, and richer attribute diversity. Some of these directions are already being explored in follow-up work [4] leveraging DisCo.
@@ E.1.3 Intra-Image Diversity Aggregation Function Analysis
E.1.4 Reinforcement Learning vs. Supervised Fine-Tuning
To isolate whether DisCo’s gains stem from RL itself rather than simply additional domain-specific training, we compare against an online SFT baseline: for each group of $M$ generated samples, the highest-reward sample is selected as a supervised target, under the same curriculum schedule and prompt dataset.
Table E.4: RL vs. online SFT on Flux-Dev. DisCo outperforms the SFT baseline on identity diversity and perceptual quality, while converging significantly faster.
Model Unique Face Accuracy (UFA) HPS Score Training Time (hrs)
Online SFT 73.0 31.0 33
DisCo (RL) 98.6 33.4 13
As shown in Table E.4, DisCo substantially outperforms online SFT on UFA (98.6% vs. 73.0%) and HPSv2 (33.4 vs. 31.0), while converging more than $2\times$ faster (13 vs. 33 hours). This gap arises from two factors: RL’s exploration and group-normalized advantage estimation encourages diverse generation strategies rather than collapsing to the mode of high-reward samples, and the group-wise diversity reward (Eq. 6) exploits GRPO’s group structure to penalize cross-sample identity repetition. This is a signal that online SFT cannot capture by optimizing each sample independently.
E.1.5 Curriculum Bound Ablation
The curriculum learning strategy anneals from a simple prompt distribution to uniform sampling over the full range. A key design choice is the upper bound of the “simple” set. This is the maximum person count used during the initial curriculum phase. Table E.5 ablates this bound on the Flux-Krea baseline, keeping all other hyperparameters fixed. The no-curriculum row (2-7) corresponds to uniform sampling from the start, equivalent to the penultimate row of Table 2.
Table E.5: Ablation on curriculum simple-set bounds on Flux-Krea baseline. Blue row is the selected configuration.
Curriculum Bound Count Accuracy Unique Face Accuracy (UFA) Global Identity Spread (GIS) HPS Score
2-7 79.2 82.6 73.7 32.4
2-3 84.3 83.0 85.5 32.3
2-4 83.5 89.7 90.6 32.2
2-6 80.1 83.1 77.0 32.2
As shown in Table E.5, no curriculum (2-7 uniform) leads to the weakest UFA and GIS, confirming that gradual complexity introduction is critical for the specialist Flux-Krea model. Starting with too narrow a range (2-3) yields strong count accuracy but leaves UFA and GIS below our chosen bound, suggesting insufficient exposure to harder scenarios during the curriculum phase. Starting too broad (2-6) performs similarly to no curriculum on GIS, as the model is exposed to complex prompts too early. The 2-4 bound has the best balance.
E.1.6 Final Run Reward Curves
@@ E.1.7 Computational Analysis
E.1.8 User Preference Study
Our proposed metrics based on face detection and embedding similarity are efficient and reproducible, but it is important to validate that they align with human perception. We conduct a user preference study to assess whether these metrics meaningfully reflect how humans judge identity diversity in generated images.
Setup.
We collect 22 pairs of multi-human images generated by DisCo and baseline Flux, spanning a range of person counts and prompt types. 25 participants were shown each pair and asked to select the image containing less visually similar faces, i.e., the one exhibiting more identity diversity. Participants were given no information about which model produced each image.
Results.
Figure E.2: User preference study results. Each point represents one image from a pair, plotted by its mean and max pairwise cosine face similarity. Green points are the images humans collectively judged as more diverse (lower similarity); red points are the images they judged as less diverse. The two clusters separate naturally in the 2D metric space, confirming that mean and max cosine similarity jointly align with human perception of identity diversity.
As shown in Figure E.2, the images humans chose as more diverse (green) and less diverse (red) form two naturally separated clusters in the space of mean and max pairwise cosine similarity. Quantitatively, our UFA metric achieves a correlation of 1.0 with human judgments, and mean pairwise cosine similarity achieves a correlation of 0.91, confirming that both the automated metrics and rewards capture what humans actually perceive as identity diversity rather than purely reflecting the behavior of the underlying face recognition model.
E.2 Qualitative Results
E.2.1 Visualizing Global Identity Spread
@@
Figure E.3: DisCo v/s Flux-Dev As observed in this Figure, we visualize three prompts of people containing the same ethnicity, over six consistent seeds for DisCo and Flux. As observed, Flux results not only generate overlapping identities in the same image, but generate similar looking people across the dataset. However, DisCo finetuning pushes the faces further from each other.
E.2.2 Results on Flux-Krea
@@
E.2.3 Visual Effects of Count and HPS Reward Components
@@
Appendix F Generalizability and Applicability of DisCo
F.1 Out-of-Distribution Evaluation for number of People
Our main experiments evaluate on prompts with 2-7 people, matching the training distribution. Here we examine whether DisCo generalizes to out-of-distribution person counts, evaluating on prompts with up to 10 individuals across 270 samples.
Figure F.1: Performance on 2-10 people. DisCo(Flux) retains high Unique Face Accuracy, Count Accuracy and HPSv2 beyond the training range of 7 people, continuing to outperform baseline Flux at 8, 9, and 10 individuals.
Figure F.1 shows that DisCo generalizes gracefully beyond its training range. Unique Face Accuracy remains high at 8-10 people, where baseline Flux degrades sharply. This is due to the fact that the identity collapse problem becomes significantly worse as group size grows, yet DisCo continues to produce distinct individuals. Count Accuracy and HPSv2 scores do tend to drop-off for both methods. Regardless, DisCo remains to significantly outperform the baseline at higher person counts. These observations demonstrate that the gains observed within the training range are not simply a product of distribution memorization but reflect a genuine improvement in the model’s ability to handle complex multi-human scenes.
F.2 Beyond Facial Diversity
DisCo is trained with facial identity diversity rewards. In this section, we study whether the learned behavior generalizes to other visual attributes that distinguish individuals in a scene. We investigate two such attributes: clothing diversity and body-type diversity.
Metrics.
Clothing diversity is measured as the average pairwise CLIP [26] similarity between clothing regions across people within each generated image. Clothing regions are obtained by cropping the torso area of each detected person. Lower similarity indicates more diverse clothing across individuals. Body-type diversity is measured as the ratio of people sharing the same somatotype within an image, assessed using Qwen3-VL [1] to classify each person’s body type from the generated image. A lower ratio indicates greater variety in body types across the generated group.
For both metrics, lower scores indicate better diversity, denoted by $\downarrow$ in Table F.1.
Quantitative Results.
Table F.1: Clothing and body-type diversity. DisCo trained on facial diversity already improves both attributes; fine-tuning with a clothing-specific reward (Fashion-SigLIP) yields further gains.
[…]
Flux-Dev 0.73 0.55
+DisCo(Face) 0.65 0.49
+DisCo(Cloth) 0.60 0.49
Table F.1 shows two findings. First, DisCo trained solely on facial diversity already improves both clothing and body-type diversity over the Flux-Dev baseline. Note that this is despite never explicitly optimizing for either attribute. We hypothesize this is because the model learns a broader notion of identity distinctiveness: generating people whose faces look different from each other naturally induces variation in clothing and body type along with the facial features. Second, replacing the ArcFace identity reward with a Fashion-SigLIP [^9] https://huggingface.co/Marqo/marqo-fashionSigLIP clothing similarity signal yields further gains in clothing diversity while retaining the body-type improvement, demonstrating that DisCo’s reward framework generalizes cleanly to other embedding spaces and attribute types beyond faces.
Qualitative results.
Figure F.2: Generalization to clothing diversity. Qualitative comparison between Flux-Dev and DisCo(Cloth). Fine-tuning with a Fashion-SigLIP clothing reward produces noticeably distinct clothes across all generated people.
Figure F.2 illustrates these trends qualitatively. The baseline Flux model tends to dress generated individuals in similar or near-identical outfits. DisCo visibly increases clothing variation as a byproduct of identity diversity training, producing groups where each person’s clothing is clearly distinct in style, color, and type.
F.3 General Text-to-Image Performance on GenEval2
A valid concern with any RL fine-tuning approach is whether optimizing for a specific objective degrades the base model’s general capabilities. We evaluate DisCo on GenEval2 [12], a comprehensive benchmark that tests compositional text-to-image generation across object presence, attributes, counting, spatial positioning, and verb understanding, as well as overall prompt complexity (atomicity 3-7+).
Table F.2: GenEval2 results. DisCo improves over baseline Flux across most compositional attributes and prompt complexity levels, demonstrating that diversity-focused RL fine-tuning does not hurt general text-to-image capability.
Attribute Type Atomicity
Object Attribute Count Position Verb 3 4 5 6 7+ Mean
Flux-Dev 89.2 67.2 53.9 41.5 19.0 61.4 39.6 27.5 29.2 8.7 24.6
+DisCo 94.5 66.9 58.7 45.1 27.2 61.2 49.2 32.2 29.9 11.0 27.6
As shown in Table F.2, DisCo improves over baseline Flux on nearly all axes, including object, counting, positioning, and verb. The mean score rises from 24.6 to 27.6. Gains are particularly notable at higher prompt complexity levels (atomicity 4-7+), which correspond to the multi-human scenarios DisCo is trained on. The only marginal regression is on attribute following (67.2 to 66.9), which we consider negligible. These results indicate that DisCo’s RL fine-tuning using count and HPS rewards not only preserves but actively improves the base model’s general compositional capabilities. This is consistent with RL’s known ability to correct ingrained model behaviors while retaining broad generalization.
F.4 Compositional Control with DisCo
DisCo is trained for identity diversity, yet we observe that fine-grained compositional control is preserved and in many cases improved. This is due to the HPSv3 reward component. Figure F.3 illustrates this: prompts with explicit per-person clothing and hairstyle specifications are faithfully followed across all individuals simultaneously.
Figure F.3: DisCo follows per-person compositional prompts. Despite being trained purely for identity diversity, DisCo successfully generates distinct individuals with specified per-person clothing and hairstyles.
Quantitative Results.
Table 1 corroborates this qualitatively. On both DiverseHumans-Test and MultiHuman-TestBench, which explicitly test complex compositional scenes with per-person action specifications, DisCo retains and slightly improves Action scores over the base model, while also improving HPSv2. This is the harder test of compositional following, as prompts describe distinct actions per individual rather than a shared scene context. We attribute this to the HPSv3 reward, which directly optimizes prompt alignment during RL fine-tuning and prevents quality regression which might otherwise arise from diversity-focused optimization. This is further supported by our GenEval2 results (Table F.2), where DisCo improves over baseline Flux on compositional attributes including count, position, and verb understanding. Future work such as Ar2Can [4] builds on this foundation toward more explicit per-person spatial and identity-level control.
Appendix G Limitations and Future Work
DisCo relies on face detection and embedding similarity, so performance can degrade under heavy occlusion, extreme poses, or very small faces. While DisCo does not explicitly train for per-person compositional attributes such as clothing or body type, we show in Appendix F.4 that the HPSv3 reward preserves and often improves the base model’s ability to follow such specifications. DisCo is not a personalization method and only text-to-image generation. However, personalization using explicit spatial and identity-level control is explored in concurrent work Ar2Can [4], which builds directly on DisCo’s RL formulation. Other future directions include extending diversity optimization to whole-body appearance cues, video generation with spatiotemporal identity consistency, and more explicit fairness and demographic balance objectives.
Appendix H Ethics Statement